# Adaptive Reconstruction for Graph Neural Networks

### v2

@@ Abstract
Graph Neural Networks (GNNs) have become fundamental in semi-supervised learning for graph representation, leveraging their ability to capture complex node relationships. A recent trend in GNN research focuses on adaptive k-hop structure learning, moving beyond fixed-hop aggregation to more flexible and dynamic neighborhood selection. While GAMLP Zhang et al. (2022) employs separate MLP layers for each k-hop domain and ImprovingTE Yao et al. (2023) enhances this by injecting contextualized substructure information, these methods still rely heavily on predefined sampling strategies, which may limit their ability to generalize and maintain stable accuracy. To address these limitations, we propose an adaptive reconstruction framework that dynamically refines k-hop structure learning. Inspired by ”coreset selection” Guo et al. (2022), our approach adaptively reconstructs node neighborhoods to optimize message passing, ensuring more effective and context-aware information flow across the graph. To further enhance structural robustness, we introduce two key modules: the Distance Recomputator and the Topology Reconstructor (DRTR). The Distance Recomputator reassesses and recalibrates node distances based on adaptive graph properties, leading to improved node embeddings that better reflect latent relationships. Meanwhile, the Topology Reconstructor dynamically refines local graph structures, enabling the model to adapt to evolving graph topologies and mitigate the impact of noise and mislabeled data. Empirical evaluations demonstrate that our adaptive reconstruction framework achieves significant improvements over existing k-hop-based models, providing more stable and accurate performance in various graph learning benchmarks.
@@ 1 Introduction
Existing GNN architectures predominantly adopt static node representations and predefined aggregation schemes, which can be suboptimal when handling graphs with dynamic topologies. Furthermore, the computation of node distances in large-scale graphs introduces significant overhead, making it challenging to scale these models effectively Kung et al. (2024); Wang et al. (2024). These limitations highlight the need for a more flexible and efficient approach that can dynamically reconstruct graph structures, allowing models to better adapt to changes in connectivity patterns and improve representation quality.
To address these challenges, we propose an adaptive reconstruction framework that refines both node distance calculations and local topological structures to enhance message passing efficiency. This framework introduces two key components: the Distance Recomputator and the Topology Reconstructor (DRTR). The Distance Recomputator dynamically recalibrates node distances within the k-hop domain by incorporating an adaptive encoding mechanism that adjusts to changes in graph density and node proximity. By refining distance computations, it provides a more precise representation of node relationships, which improves downstream predictions and model stability. The Topology Reconstructor complements this by continuously adjusting local graph connectivity, ensuring that the model remains robust in scenarios where structural variations occur. This adaptive mechanism allows GNNs to dynamically reconstruct neighborhood structures, making them more resilient to graph shifts and variations in node interactions.
@@ 2.1 Topology Imbalance Problem
The structural connections in a graph do not always accurately reflect the significance of each link. As stated in Liu et al. (2023), ”not every link is as useful as its topology suggests.” Consider a social network where an account labeled ”math” follows ten mathematical educators and one food-maker. For tasks such as label prediction, the connection to the food-maker contributes little relevant information, yet traditional Graph Neural Networks (GNNs) aggregate information from all 1-hop neighbors indiscriminately. This results in a phenomenon we define as the Topology Imbalance Problem, where some connections (e.g., links to math educators) carry significantly more useful information than others (e.g., the food-maker link).
@@ 2.2 Limitations of Synchronous Aggregation
To address this limitation, we propose an asynchronous aggregation mechanism that selectively integrates information from different neighbors based on their relevance. Unlike traditional synchronous approaches, our method dynamically determines the order and priority of aggregation, reducing the negative influence of uninformative links while preserving critical structural relationships. Some recent works have explored asynchronous processing in GNNs, such as AEGNN Schaefer et al. (2022), which updates node activations only upon new events, and Gated Graph Sequence Neural Networks Li et al. (2017), which employ gated recurrent units to process sequences asynchronously. However, these methods primarily focus on message passing rather than addressing topological reconstruction at a structural level. Our approach directly enhances the representation of graph topology, offering a more comprehensive solution.
@@ 2.3 Heat Diffusion in Graph Neural Networks
The concept of heat diffusion offers a valuable perspective for understanding information propagation in Graph Neural Networks (GNNs). Drawing inspiration from thermodynamics, where heat gradually dissipates as it moves away from its source Thanou et al. (2016), we analyze how information in GNNs follows a similar pattern of attenuation during multi-hop message passing. As information propagates across a graph, its influence weakens with increasing distance from the source node, often leading to a decline in signal quality and an accumulation of noise. Traditional GNN architectures typically assume uniform propagation across all hops, failing to account for this natural decay in information relevance.
@@ 2.4 Paper Contribution
  • Distance Recomputator: Our model recalculates node distances using an adaptive attention mechanism that considers both k-hop topology and feature interactions. It continuously adjusts distances based on contextual relevance, improving message passing and filtering irrelevant edges.
  • Topology Reconstructor: We introduce a k-hop topology reconstruction method. our model leverages k-hop topology from sampling to perform reconstruction based on computed ”similarity distances.” Nodes exceeding a similarity distance threshold are repositioned to optimize network configurations for enhanced learning outcomes.
@@ 3.1 K-hop Message Passing
Most existing k-hop models in graph neural networks rely on multiple rounds of 1-hop message passing to capture higher-order information. However, they lack the ability to perform k-hop message passing with k-hop propagation in a single iteration or efficiently sample k-hop neighborhoods in one step. For example, the enhancing multi-hop connectivity model refines k-hop neighborhoods by sampling k-hop neighbors and reweighting high-order connections, strengthening highly related nodes while down-weighting weaker ones Liu et al. (2022). Similarly, KP-GNN enriches node representations by incorporating peripheral embeddings at each layer Feng et al. (2022).
Our model introduces an efficient k-hop sampling strategy that enables direct k-hop message passing, capturing local structures more effectively than traditional 1-hop methods. By integrating k-hop information in each propagation step, our approach provides a more comprehensive graph perspective, leading to improved label prediction accuracy for central nodes.
@@ 3.2 Diffusion Propagation in Graph Neural Networks
We propose a diffusion-based propagation model that builds upon traditional approaches such as the Diffusion Decent Network. One well-known model, MAGNA Wang et al. (2021), propagates information by updating vertex and edge embeddings while recalculating distances for both one-hop and multi-hop neighbors, considering all paths up to $k$ hops. Another approach, ”Diffusion Improves Graph Learning” Klicpera et al. (2019), approximates the diffusion equation using an infinite series, improving computational efficiency on large-scale graphs.
Our model extends these concepts by simulating heat diffusion within the k-hop domain. This diffusion-inspired framework regulates message passing, dynamically adjusting the flow of information to reduce noise accumulation and enhance representation quality. By incorporating heat diffusion into k-hop propagation, our model provides a structured and adaptive approach to information distribution in graph neural networks.
@@ 3.3 Graph Imbalance Problems
Label and topology imbalance in graph neural networks remain significant but underexplored challenges. Zhao et al. Zhao et al. (2022) addressed topology imbalance by adjusting edge weights without modifying the graph structure.
@@ 4 Architectures of DRTR
In GNN tasks, the topology information and node feature are combined to learn the representation vetor of a node for node tasks. Modern GNNs aggregate the information from nodes in neighborhood and update the representation of the nodes by a message-passing scheme. After k-iterations of aggregation, the representation of a node captures the structural information within its k-hop neighborhood. Formally, the layer-wise aggregation of a GNN is given by
@@ 4.1 In-Depth Learning in Graph Neural Networks
Contemporary GNNs predominantly employ a 1-hop message passing paradigm, where each propagation step explores information within a single depth level. Unlike GraphSAGE Hamilton et al. (2017), which samples along a $k$-depth domain but aggregates at every depth, our approach integrates both sampling and aggregation within a $k$-hop neighborhood. This enables a more comprehensive understanding of local structural dependencies within the graph.
@@
The $k$-hop attentive message passing strategy leverages both static and dynamic sampling pathways. This mechanism is inspired by graph diffusion networks and integrates distance reevaluation indicators derived from the Graph Attention Network (GAT) Veličković et al. (2018). By synthesizing static and dynamic sampling methodologies, our approach enhances both the depth and accuracy of message passing in GNNs.
@@
DRTR Diffusion Layer design is an integration of K-hop diffusion, adaptive message passing, and dynamic topology reconstruction to capture both local and long-range dependencies. Instead of standard 1-hop aggregation, DRTR employs an iterative K-hop aggregation process, where node representations are refined using a combination of local neighborhood aggregation and global diffusion attention. To further improve feature propagation, a Graph Attention Mechanism (GAT) dynamically computes attention weights, prioritizing important neighbors and filtering out noisy connections. Additionally, DRTR introduces a Topology Reconstructor Module, which adjusts edge weights and graph connectivity by recomputing distances and adapting to evolving structures. This mechanism mitigates over-smoothing and ensures the model remains responsive to structural variations. By balancing computational efficiency and expressive power, DRTR’s hierarchical multi-hop framework enables effective large-scale graph learning while maintaining a trade-off between depth (global context) and locality (fine-grained details).
@@
In this paper, we adopt a structure that considers both k-hop sampling and k-hop message passing within the graph diffusion network framework. Striking a balance between computational efficiency (time) and model performance, our approach involves static preprocessing for initial k-hop neighborhood sampling, followed by dynamic resampling in subsequent iterations. This methodology ensures that our model remains efficient while effectively capturing the complex dependencies in the graph structure across multiple hops.
@@ 4.2 Distance Recomputator and Topology Reconstructor Implementation
In summary, our algorithm presents a novel framework that combines k-hop sampling with dynamic resampling and recomputation, all underpinned by an asynchronous aggregation strategy. This comprehensive approach addresses key challenges in GNNs, such as handling complex graph topologies and efficiently updating node representations, thereby setting a new standard in the field of graph neural network research.
@@ 4.3 Experimental Analysis
  • Citeseer: For Citeseer, the SSGC+GKHDDRA combination achieved the highest accuracy at 75.6%. This underscores the effectiveness of incorporating k-hop sampling and dynamic resampling in dealing with the dataset’s complex graph structure.

### v3

# Adaptive Topology Reconstruction for Robust Graph Representation Learning

@@ Abstract
Graph Neural Networks (GNNs) have become fundamental in semi-supervised learning for graph representation, leveraging their ability to capture complex node relationships. A recent trend in GNN research focuses on adaptive multi-hop structure learning, moving beyond fixed-hop aggregation to more flexible and dynamic neighborhood selection. While GAMLP (Zhang et al., 2022) employs separate MLP layers for each multi-hop domain and ImprovingTE (Yao et al., 2023) enhances this by injecting contextualized substructure information, these methods still rely heavily on predefined sampling strategies, which may limit their ability to generalize and maintain stable accuracy. To address these limitations, we propose an adaptive reconstruction framework that dynamically refines multi-hop structure learning. Inspired by ”coreset selection” (Guo et al., 2022), our approach adaptively reconstructs node neighborhoods to optimize message passing, ensuring more effective and context-aware information flow across the graph. To further enhance structural robustness, we introduce two key modules: the Distance Recomputator and the Topology Reconstructor (DRTR). The Distance Recomputator reassesses and recalibrates node distances based on adaptive graph properties, leading to improved node embeddings that better reflect latent relationships. Meanwhile, the Topology Reconstructor dynamically refines local graph structures, enabling the model to adapt to evolving graph topologies and mitigate the impact of noise and mislabeled data. Empirical evaluations demonstrate that our adaptive reconstruction framework achieves significant improvements over existing multi-hop-based models, providing more stable and accurate performance in various graph learning benchmarks.
@@ 1 Introduction
Existing GNN architectures predominantly adopt static node representations and predefined aggregation schemes, which can be suboptimal when handling graphs with dynamic topologies. Furthermore, the computation of node distances in large-scale graphs introduces significant overhead, making it challenging to scale these models effectively (Kung et al., 2024; Wang et al., 2024). These limitations highlight the need for a more flexible and efficient approach that can dynamically reconstruct graph structures, allowing models to better adapt to changes in connectivity patterns and improve representation quality.
To address these challenges, we propose an adaptive reconstruction framework that refines both node distance calculations and local topological structures to enhance message passing efficiency. This framework introduces two key components: the Distance Recomputator and the Topology Reconstructor (DRTR). The Distance Recomputator dynamically recalibrates node distances within the multi-hop domain by incorporating an adaptive encoding mechanism that adjusts to changes in graph density and node proximity. By refining distance computations, it provides a more precise representation of node relationships, which improves downstream predictions and model stability. The Topology Reconstructor complements this by continuously adjusting local graph connectivity, ensuring that the model remains robust in scenarios where structural variations occur. This adaptive mechanism allows GNNs to dynamically reconstruct neighborhood structures, making them more resilient to graph shifts and variations in node interactions.
@@ 2.1 Topology Imbalance Problem
The structural connections in a graph do not always accurately reflect the significance of each link. As stated in (Liu et al., 2023), ”not every link is as useful as its topology suggests.” Consider a social network where an account labeled ”math” follows ten mathematical educators and one food-maker. For tasks such as label prediction, the connection to the food-maker contributes little relevant information, yet traditional Graph Neural Networks (GNNs) aggregate information from all 1-hop neighbors indiscriminately. This results in a phenomenon we define as the Topology Imbalance Problem, where some connections (e.g., links to math educators) carry significantly more useful information than others (e.g., the food-maker link).
@@ 2.2 Limitations of Synchronous Aggregation
To address this limitation, we propose an asynchronous aggregation mechanism that selectively integrates information from different neighbors based on their relevance. Unlike traditional synchronous approaches, our method dynamically determines the order and priority of aggregation, reducing the negative influence of uninformative links while preserving critical structural relationships. Some recent works have explored asynchronous processing in GNNs, such as AEGNN (Schaefer et al., 2022), which updates node activations only upon new events, and Gated Graph Sequence Neural Networks (Li et al., 2017), which employ gated recurrent units to process sequences asynchronously. However, these methods primarily focus on message passing rather than addressing topological reconstruction at a structural level. Our approach directly enhances the representation of graph topology, offering a more comprehensive solution.
@@ 2.3 Heat Diffusion in Graph Neural Networks
The concept of heat diffusion offers a valuable perspective for understanding information propagation in Graph Neural Networks (GNNs). Drawing inspiration from thermodynamics, where heat gradually dissipates as it moves away from its source (Thanou et al., 2016), we analyze how information in GNNs follows a similar pattern of attenuation during multi-hop message passing. As information propagates across a graph, its influence weakens with increasing distance from the source node, often leading to a decline in signal quality and an accumulation of noise. Traditional GNN architectures typically assume uniform propagation across all hops, failing to account for this natural decay in information relevance.
@@ 2.4 Paper Contribution
  • Distance Recomputator: Our model recalculates node distances using an adaptive attention mechanism that considers both multi-hop topology and feature interactions. It continuously adjusts distances based on contextual relevance, improving message passing and filtering irrelevant edges.
  • Topology Reconstructor: We introduce a multi-hop topology reconstruction method. our model leverages multi-hop topology from sampling to perform reconstruction based on computed ”similarity distances.” Nodes exceeding a similarity distance threshold are repositioned to optimize network configurations for enhanced learning outcomes.
@@ 3.1 K-hop Message Passing
Most existing multi-hop models in graph neural networks rely on multiple rounds of 1-hop message passing to capture higher-order information. However, they lack the ability to perform k-hop message passing with multi-hop propagation in a single iteration or efficiently sample multi-hop neighborhoods in one step. For example, the enhancing multi-hop connectivity model refines multi-hop neighborhoods by sampling multi-hop neighbors and reweighting high-order connections, strengthening highly related nodes while down-weighting weaker ones (Liu et al., 2022). Similarly, KP-GNN enriches node representations by incorporating peripheral embeddings at each layer (Feng et al., 2022).
Our model introduces an efficient multi-hop sampling strategy that enables direct k-hop message passing, capturing local structures more effectively than traditional 1-hop methods. By integrating multi-hop information in each propagation step, our approach provides a more comprehensive graph perspective, leading to improved label prediction accuracy for central nodes.
@@ 3.2 Diffusion Propagation in Graph Neural Networks
We propose a diffusion-based propagation model that builds upon traditional approaches such as the Diffusion Decent Network. One well-known model, MAGNA (Wang et al., 2021), propagates information by updating vertex and edge embeddings while recalculating distances for both one-hop and multi-hop neighbors, considering all paths up to $k$ hops. Another approach, ”Diffusion Improves Graph Learning” (Klicpera et al., 2019), approximates the diffusion equation using an infinite series, improving computational efficiency on large-scale graphs.
Our model extends these concepts by simulating heat diffusion within the multi-hop domain. This diffusion-inspired framework regulates message passing, dynamically adjusting the flow of information to reduce noise accumulation and enhance representation quality. By incorporating heat diffusion into multi-hop propagation, our model provides a structured and adaptive approach to information distribution in graph neural networks.
@@ 3.3 Graph Imbalance Problems
Label and topology imbalance in graph neural networks remain significant but underexplored challenges. Zhao et al. (Zhao et al., 2022) addressed topology imbalance by adjusting edge weights without modifying the graph structure.
@@ 4 Architectures of DRTR
In GNN tasks, the topology information and node feature are combined to learn the representation vetor of a node for node tasks. Modern GNNs aggregate the information from nodes in neighborhood and update the representation of the nodes by a message-passing scheme. After k-iterations of aggregation, the representation of a node captures the structural information within its multi-hop neighborhood. Formally, the layer-wise aggregation of a GNN is given by
@@ 4.1 In-Depth Learning in Graph Neural Networks
Contemporary GNNs predominantly employ a 1-hop message passing paradigm, where each propagation step explores information within a single depth level. Unlike GraphSAGE (Hamilton et al., 2017), which samples along a $k$-depth domain but aggregates at every depth, our approach integrates both sampling and aggregation within a $k$-hop neighborhood. This enables a more comprehensive understanding of local structural dependencies within the graph.
@@
The $k$-hop attentive message passing strategy leverages both static and dynamic sampling pathways. This mechanism is inspired by graph diffusion networks and integrates distance reevaluation indicators derived from the Graph Attention Network (GAT) (Veličković et al., 2018). By synthesizing static and dynamic sampling methodologies, our approach enhances both the depth and accuracy of message passing in GNNs.
@@
DRTR Diffusion Layer design is an integration of multi-hop diffusion, adaptive message passing, and dynamic topology reconstruction to capture both local and long-range dependencies. Instead of standard 1-hop aggregation, DRTR employs an iterative K-hop aggregation process, where node representations are refined using a combination of local neighborhood aggregation and global diffusion attention. To further improve feature propagation, a Graph Attention Mechanism (GAT) dynamically computes attention weights, prioritizing important neighbors and filtering out noisy connections. Additionally, DRTR introduces a Topology Reconstructor Module, which adjusts edge weights and graph connectivity by recomputing distances and adapting to evolving structures. This mechanism mitigates over-smoothing and ensures the model remains responsive to structural variations. By balancing computational efficiency and expressive power, DRTR’s hierarchical multi-hop framework enables effective large-scale graph learning while maintaining a trade-off between depth (global context) and locality (fine-grained details).
@@
In this paper, we adopt a structure that considers both k-hop sampling and k-hop message passing within the graph diffusion network framework. Striking a balance between computational efficiency (time) and model performance, our approach involves static preprocessing for initial multi-hop neighborhood sampling, followed by dynamic resampling in subsequent iterations. This methodology ensures that our model remains efficient while effectively capturing the complex dependencies in the graph structure across multiple hops.
@@ 4.2 Distance Recomputator and Topology Reconstructor Implementation
In summary, our algorithm presents a novel framework that combines multi-hop sampling with dynamic resampling and recomputation, all underpinned by an asynchronous aggregation strategy. This comprehensive approach addresses key challenges in GNNs, such as handling complex graph topologies and efficiently updating node representations, thereby setting a new standard in the field of graph neural network research.
@@ 4.3 Experimental Analysis
  • Citeseer: For Citeseer, the SSGC+GKHDDRA combination achieved the highest accuracy at 75.6%. This underscores the effectiveness of incorporating multi-hop sampling and dynamic resampling in dealing with the dataset’s complex graph structure.